Hi all I need some help with this one please. I am not sure of the result I find.

You are working on your second project as an equity research intern at a bulge investment bank. Your focus is in retail space, especially in the health and fitness sector. Currently, you are gathering information on a fast-growing chain fitness company called LuluYoga. You are interested in calculating the free cash flow of the firm.

LuluYoga offers yoga classes in several major cities in the United States. Two major revenue resources are selling workout gear and membership passes for class access.

Assume at the beginning of year 2016, LuluYoga has zero inventory.

In year 2016, LuluYoga purchased 10,000 yoga mats at a price of $10 each. The company sells 6,000 mats at a price of $15 in year 2016 and sells the remaining at a price of $20 in year 2017.

In year 2016, LuluYoga sells 1,000 membership passes for $2,000 each. 80% of the classes purchased were used in 2016 and the rest are used in 2017.The yoga master’s compensation to teach classes are $300K in year 2016 and $200K in year 2017.

LuluYoga pays corporate tax of 35%.

Question is :

What is the FCF of LuluYoga in year 2016 and in year 2017
Answer #1

Free Cash Flow of the Firm (FCFF)

FCFF can be understood from the below formula,

FCFF= earnings before interest and taxes (EBIT) x (1 - tax rate) + depreciation - long-term investments - investments in working capital

FCFF for the year 2016

Calculation of EBIT,

Profit on sale of Yoga Mat = $(90000-60000) = $30000/-

Profit on sale of membership passes = $(1000*80%)*2000 - 300000 = $ 1300000/-

Total EBIT = $1330000/-

Working Capital = current assets - current liabilities

WC = (4000*10) - (1000*20%)*2000

WC = $-360000

FCFF = 1330000(1-35%) - 360000

FCFF = $504500

FCFF for the year 2017

Profit on sale of Yoga Mat = $4000*(15-10) = $20000/-

Profit on sale of membership passes = $(1000*20%)*2000 - 200000 = $ 200000/-

Total EBIT = $220000/-

Working Capital = NIL

FCFF = 220000*(1-35%) = $ 143000/-